I have to use a system of two linear equations to solve this exercise. The first problem is : the sum of 2 numbers is 87. If twice the smaller number is subtracted from the larger number, the result is 12. Find the two numbers. Please help. Thank you
Answer by stanbon(75887)

the sum of 2 numbers is 87. If twice the smaller number is subtracted from the larger number, the result is 12. Find the two numbers.
----------
Equations:
a + b = 87
a - 2b = 12
------------------
Subtract and solve for "b"::
3b = 75
b = 25
----
a + 25 = 87
a = 62
-------------
Cheers,
Stan H.
-------------